Solve this equation to get the absolute value of:
[(2/7x + 4/7)-3] = 5
How do you get the answers : -30 and 26?

Answer by t0hierry(194)

You know you will get two answers, depending on whether the expression is positive or negative.
if 2/7x + 4/7 - 3 is positive then the equation is 2/7x + 4/7 - 3 = 5
or 2/7x = 5 + 3 -4/7 that is 2x = 35 + 21 -4 = 52 or x = 26
When is 2/7x + 4/7 -3 = 0
When 2x = 4 - 21 ie x = -17/2
mmh Can you show me the absolute value?